# Diffusion-Driven Generative Framework for Molecular Conformation Prediction

### 22 Dec 2023 v1

@@ V Experiments
In the ensuing paragraphs, we subject DDGF to rigorous empirical testing, focusing on its efficacy in generating equilibrium conformations for molecules of various sizes, including those that are pertinent to pharmacological interests. In line with established precedents in the field [25, 30], we scrutinize our approach alongside other prominent contenders across two benchmark tests. The first, Conformation Generation, gauges the proficiency of the model in capturing the conformational distribution by assessing both the diversity and the fidelity of the conformations it generates. The second, Property Prediction, examines the model’s ability to infer molecular properties accurately. We outline the general setup for our experiments before delving into the evaluation criteria tailored to each task and the subsequent discussion of the findings.
@@ V-A1 Datasets
Building on the frameworks set by recent investigations [18], our analysis incorporates data from the GEOM-QM9 [49] and GEOM-Drugs [30] datasets, each hosting a spectrum of molecular structures. GEOM-QM9 is renowned for its collection of small molecules, whereas GEOM-Drugs encompasses medium-sized organic compounds. We adopt the dataset partitions established by Shi et al. [25], with both datasets’ training splits encompassing 40,000 molecules, each represented by 5 distinct conformations, cumulating in a total of 200,000 conformations. The validation set mirrors the size of the training split. For our evaluation, the test split presents 200 unique molecules, translating to 22,408 conformations within QM9 and 14,324 for the Drugs dataset. This rich repository of structural data sets the stage for our comprehensive model assessments.
The GEOM-QM9 dataset extends the original QM9 collection by including multiple conformations for each molecule, focusing on smaller compounds with up to 9 heavy atoms. For our training, we extracted 50,000 pairs of conformations and molecules from this dataset, and set aside 17,813 conformations from 150 molecular structures for testing. In contrast, the GEOM-Drugs dataset comprises larger molecules with an average of 6.5 rotatable bonds, indicative of complex molecular dynamics. From this, we formed our training set with another 50,000 pairs and allocated 9,161 conformations from 100 molecular structures for testing, allowing us to assess our model’s performance across a wide range of molecular complexities.
@@ V-A2 Evaluation metrics
Evaluating the conformation generation models hinges on their ability to produce diverse and accurate molecular structures, for which the root-mean-square deviation (RMSD) serves as a critical measure after aligning the structures using the Kabsch algorithm. Metrics include Coverage, which reflects the model’s diversity by indicating the proportion of reference conformations matched within a certain RMSD threshold, and Matching, which mirrors precision by showing how many generated conformations closely resemble those in the reference set. Together with the Minimum and Maximum RMSD metrics that respectively shed light on the average best-case alignment and the worst outliers, these tools collectively gauge the model’s efficacy in crafting conformations that are both varied and true to real molecular shapes.
@@
[…]
The property prediction task appraises the molecular ensemble properties over a range of generated conformations, providing a direct evaluation of sample quality. Specifically, we adopt the approach detailed in previous research where a subset of 30 molecules from the GEOM-QM9 dataset is used to generate 50 samples for each molecule. Utilizing the chemical computation toolkit PSI4, we calculate the energy ($E$) and HOMO-LUMO gap ($\epsilon$) for each conformer. These calculations allow us to compare several statistics, the average energy ($\bar{E}$), the lowest energy conformer ($E_{min}$), the average HOMO-LUMO gap ($\bar{\Delta\epsilon}$), the smallest gap ($Delta\epsilon_{min}$), and the largest gap ($\Delta\epsilon_{max}$), against the ground truth values. Discrepancies in these values serve as indicators of the generative model’s accuracy in reflecting true molecular properties.
@@ V-C Results and Analysis
The summarized results in Table I depict a comprehensive performance assessment of molecular conformation generation methods. The table is bifurcated into two sections, where the top five rows feature deep generative models, and the bottom six rows include various methods enhanced with an additional rule-based force field for optimizing the generated structures. The reported metrics include Coverage (COV) and Matching (MAT) scores, with Mean and Median values calculated across different molecular graphs within the GEOM test set. For a fair and consistent evaluation [18], the size of the generated set is double that of the reference set. The standout performance of DDGF, as reflected in the tables, indicates its exceptional capability to model the multimodal distribution of molecular conformations, delivering both precision and diversity. This is especially evident with the more complex and larger molecules found in the Drugs dataset, where DDGF surpasses state-of-the-art machine learning (ML) models by a significant margin. An interesting observation is the slight edge of DDGF-C over DDGF-A, hinting at the efficacy of the chain-rule approach in enhancing the optimization process. Given this finding, DDGF-C is selected for further comparisons.
@@
On the intricate Drugs dataset, the performance of RDKIT was also evaluated. According to the data presented in Table I, there is a consensus with previous studies that state-of-the-art machine learning (ML) models surpass RDKIT in Coverage-Recall (COV-R) and Matching-Recall (MAT-R) metrics. Nevertheless, when it comes to the newly introduced Precision-based metrics, ML models still fall short. This discrepancy suggests that while ML models are more exploratory in identifying a broader range of potential structures, RDKIT tends to focus on producing a smaller set of the most probable structures, emphasizing quality over quantity. The tendency of RDKIT to produce higher quality at the expense of diversity has been previously attributed to its use of an empirical force field (FF), a strategy employed in traditional computational chemistry to refine and optimize molecular structures. In an attempt to level the playing field and conduct a more equitable comparison, DDGF was also combined with an empirical force field. The resulting DDGF + FF model was able to maintain the superior diversity (as reflected by the Recall metrics) that ML models are known for, while also achieving a significant boost in accuracy (as evidenced by the Precision metrics). This outcome suggests that integrating DDGF with a force field is a potent strategy, ensuring that the model retains its capacity to generate a wide range of conformations (thereby capturing the complex multimodal nature of molecular structures), while also enhancing the precision and reliability of the conformations it predicts. It’s an effective fusion of the explorative power of ML with the targeted refinement capabilities of force fields.
@@ V-D Property Prediction
The mean absolute errors (MAE) in the context of molecular modeling are indicative of how accurately a model can predict certain properties of molecules, such as their energy levels, when compared to the ground truth obtained from experimental or highly accurate quantum chemical calculations. In this case, Table II reports the MAE for various properties of molecules that have been generated by different models. The exclusion of CVGAE from this comparison due to its subpar performance, as noted in previous work, suggests that it was significantly less accurate than other models tested. The properties in question are highly sensitive to the geometric structure of the molecules – for instance, slight changes in the distance between atoms can lead to significant differences in the predicted properties. Therefore, a model that can accurately predict these properties demonstrates that it can consistently generate molecular conformations that closely resemble true molecular structures. In summary, the low MAE reported for DDGF implies that it not only generates conformations with high diversity but also with a high degree of geometric accuracy, thus effectively predicting the properties that are closely aligned with what would be expected of the actual molecules. This reinforces the idea that DDGF is capable of capturing the complex underlying distribution of molecular structures in a way that is both diverse and precise.
@@ VI Conclusion
In conclusion, this study has unveiled a pioneering generative framework, DDGF, which signifies a substantial leap forward in the accurate prediction of three-dimensional molecular structures from their two-dimensional graph representations. By leveraging the principles of diffusion from non-equilibrium thermodynamics, DDGF transcends traditional bifurcated methodologies, instead fostering a coherent transformation of stochastic distributions into precise molecular conformations. Its unique bilevel optimization process, tailored to the intricacies of molecular geometry, and its commitment to preserving roto-translational invariance set new benchmarks in the field of computational chemistry. The extensive empirical evaluations presented herein underscore DDGF’s superiority over existing state-of-the-art methods, cementing its role as a critical tool for advancements in drug development and the broader understanding of molecular dynamics and interactions.

### 21 Jan 2024 v2

@@ V Experiments
In the following sections, we will subject DDGF to comprehensive empirical testing, with a particular focus on its effectiveness in generating equilibrium conformations for molecules of various sizes, including those of significance in the field of pharmacology. To ensure a thorough evaluation, we will follow established practices in the field [27, 32] and compare our approach with other prominent contenders through two benchmark tests. The first benchmark, known as Conformation Generation, will assess the model’s capability to capture the conformational distribution. This evaluation will consider both the diversity and the fidelity of the conformations generated by the model. The second benchmark, referred to as Property Prediction, will examine the model’s accuracy in predicting molecular properties. Before delving into the specific evaluation criteria for each task and discussing our findings, we will provide an overview of the general setup for our experiments.
@@ V-A1 Datasets
Our analysis is based on data from two primary datasets: GEOM-QM9 [51] and GEOM-Drugs [32]. These datasets encompass a wide range of molecular structures, with GEOM-QM9 primarily containing small molecules and GEOM-Drugs including medium-sized organic compounds. We have adopted the dataset partitions as defined by [27]. Here is an overview of the dataset sizes and partitions:
  • Training splits for both datasets consist of 40,000 molecules, each represented by 5 distinct conformations, resulting in a total of 200,000 conformations for training.
  • Validation sets are of the same size as the training splits.
  • The test split consists of 200 unique molecules, comprising 22,408 conformations for the QM9 dataset and 14,324 for the Drugs dataset.
This extensive collection of structural data provides a robust foundation for our comprehensive model evaluations.
The GEOM-QM9 dataset is an extension of the original QM9 dataset, enriched with multiple conformations for each molecule, primarily focusing on smaller compounds with a maximum of 9 heavy atoms. Our training data was obtained by extracting 50,000 pairs of conformations and molecules from this dataset. For testing purposes, we reserved 17,813 conformations derived from 150 molecular structures. In contrast, the GEOM-Drugs dataset contains larger molecules characterized by an average of 6.5 rotatable bonds, reflecting complex molecular dynamics. From this dataset, we constructed our training set using an additional 50,000 pairs of conformations and molecules. For evaluation, we set aside 9,161 conformations originating from 100 molecular structures. This setup enables us to assess the performance of our model across a broad spectrum of molecular complexities.
@@ V-A2 Evaluation metrics
Evaluating the performance of conformation generation models involves assessing their ability to produce a diverse set of accurate molecular structures. The root-mean-square deviation (RMSD) metric is a crucial measure in this context, which quantifies the structural differences between the generated conformations and a reference set after aligning them using algorithms like Kabsch. To comprehensively evaluate the models, several metrics are employed, including:
  • Coverage: This metric measures the diversity of the generated conformations by indicating the proportion of reference conformations that are matched within a specified RMSD threshold.
  • Matching: Matching reflects the precision of the model by quantifying how many generated conformations closely resemble those in the reference set within a defined RMSD threshold.
  • Minimum RMSD: This metric provides insight into the average best-case alignment between the generated conformations and the reference set, indicating the overall accuracy of the model.
  • Maximum RMSD: The maximum RMSD highlights the worst outliers among the generated conformations, revealing cases where the model may struggle to produce accurate structures.
These metrics collectively offer a comprehensive assessment of the model’s ability to generate diverse and accurate molecular conformations.
@@
The COV-P and MAT-P metrics provide a different perspective on evaluating conformation generation models by reversing the roles of the generated and reference sets. These metrics assess the performance of the models by doubling the size of the generated set $S_{g}$ relative to the reference set $S_{r}$ for each molecule. Here’s how they work:
  • COV-P (Coverage-Precision): COV-P rates represent the percentage of reference structures that are adequately represented within the generated set based on a pre-defined RMSD threshold $\delta$. It measures how well the generated set covers the reference structures within the specified RMSD threshold. High COV-P rates indicate a good coverage of reference structures by the generated set.
  • MAT-P (Mean RMSD-Precision): MAT-P scores reflect the mean RMSD between each generated conformation and its nearest reference counterpart. It calculates the average structural deviation between the generated and reference conformations. Low MAT-P scores indicate that the generated conformations closely resemble the reference structures in terms of structural similarity.
Indeed, these metrics, including Precision, Recall, COV-P, and MAT-P, offer a well-rounded evaluation of a conformation generation model’s performance. The selection of the RMSD threshold $\delta$ is crucial, as it determines what level of structural similarity is considered acceptable. Setting it to 0.5 Ångströms for the QM9 dataset and 1.25 Ångströms for the Drugs dataset is a common practice and aligns with established standards in the field. These threshold values are chosen to balance the need for accuracy and diversity in generated molecular structures, taking into account the specific characteristics of each dataset and the complexity of the molecules involved.
The property prediction task is essential for evaluating how well a generative model captures molecular properties across a range of generated conformations. It provides a direct assessment of the quality of the generated samples. In this task, a subset of 30 molecules from the GEOM-QM9 dataset is used to generate 50 samples for each molecule. The properties of interest, such as energy ($E$) and the HOMO-LUMO gap ($\epsilon$), are calculated for each generated conformation using the PSI4 chemical computation toolkit. The following statistics are then compared against ground truth values to assess the model’s performance:
  • The average energy ($\bar{E}$);
  • The lowest energy conformer ($E_{min}$);
  • The average HOMO-LUMO gap ($\bar{\Delta\epsilon}$);
  • The smallest gap ($\Delta\epsilon_{min}$);
  • The largest gap ($\Delta\epsilon_{max}$).
Discrepancies between the predicted values and ground truth values for these properties serve as indicators of the generative model’s accuracy in reflecting true molecular properties. This evaluation provides valuable insights into how well the model captures the energetics and electronic properties of generated molecular conformations.
@@ V-C Results and Analysis
The summarized results in Table I provide a comprehensive performance assessment of molecular conformation generation methods. The table is divided into two sections, with the top five rows featuring deep generative models and the bottom six rows including various methods enhanced with an additional rule-based force field for optimizing the generated structures. The reported metrics include Coverage (COV) and Matching (MAT) scores, with Mean and Median values calculated across different molecular graphs within the GEOM test set. To ensure a fair and consistent evaluation, the size of the generated set is double that of the reference set, following the approach used in previous research [19].
The standout performance of DDGF, as reflected in the tables, indicates its exceptional capability to model the multimodal distribution of molecular conformations, delivering both precision and diversity. This is especially evident with the more complex and larger molecules found in the Drugs dataset, where DDGF surpasses state-of-the-art machine learning (ML) models by a significant margin. An interesting observation is the slight edge of DDGF-C over DDGF-A, hinting at the efficacy of the chain-rule approach in enhancing the optimization process. Given this finding, DDGF-C is selected for further comparisons. These results demonstrate the effectiveness of DDGF in generating diverse and accurate molecular conformations, making it a promising tool for various applications in computational chemistry and drug discovery.
@@
The performance of RDKIT on the intricate Drugs dataset was evaluated, and the results are presented in Table I. These results show a consensus with previous studies that state-of-the-art machine learning (ML) models surpass RDKIT in Coverage-Recall (COV-R) and Matching-Recall (MAT-R) metrics. However, ML models still fall short when it comes to the newly introduced Precision-based metrics. This discrepancy suggests that while ML models are more exploratory in identifying a broader range of potential structures, RDKIT tends to focus on producing a smaller set of the most probable structures, emphasizing quality over quantity. This tendency of RDKIT to produce higher quality at the expense of diversity has been attributed to its use of an empirical force field (FF), a strategy employed in traditional computational chemistry to refine and optimize molecular structures. In an attempt to level the playing field and conduct a more equitable comparison, DDGF was combined with an empirical force field to create the DDGF + FF model. The results show that the DDGF + FF model was able to maintain the superior diversity (as reflected by the Recall metrics) that ML models are known for, while also achieving a significant boost in accuracy (as evidenced by the Precision metrics). This outcome suggests that integrating DDGF with a force field is a potent strategy, ensuring that the model retains its capacity to generate a wide range of conformations (thereby capturing the complex multimodal nature of molecular structures), while also enhancing the precision and reliability of the conformations it predicts. It’s an effective fusion of the exploratory power of ML with the targeted refinement capabilities of force fields.
@@ V-D Property Prediction
The mean absolute errors (MAE) reported in Table II for various properties of molecules generated by different models provide insights into how accurately a model can predict these properties when compared to ground truth data obtained from experimental or highly accurate quantum chemical calculations. The properties considered in this evaluation are highly sensitive to the geometric structure of molecules, meaning that even slight changes in the distances between atoms can lead to significant differences in the predicted properties. Therefore, a model with low MAE for these properties demonstrates its ability to consistently generate molecular conformations that closely resemble true molecular structures. In summary, the low MAE reported for DDGF suggests that it not only generates conformations with high diversity but also with a high degree of geometric accuracy. This accuracy translates into effective predictions of properties that are closely aligned with the actual properties of the molecules. This reinforces the idea that DDGF is capable of capturing the complex underlying distribution of molecular structures in a way that is both diverse and precise.
@@ VI Conclusion
In conclusion, this study introduces a groundbreaking generative framework, DDGF, that represents a significant advancement in the accurate prediction of three-dimensional molecular structures from their two-dimensional graph representations. By drawing inspiration from non-equilibrium thermodynamics and leveraging diffusion principles, DDGF offers a cohesive approach to transforming stochastic distributions into precise molecular conformations. Its novel bilevel optimization process, designed to address the complexities of molecular geometry, and its commitment to maintaining roto-translational invariance set new standards in the field of computational chemistry. The extensive empirical evaluations presented in this study demonstrate DDGF’s superiority over existing state-of-the-art methods, establishing its significance as a vital tool for advancements in drug development and a deeper understanding of molecular dynamics and interactions.